Find the zeros of the function.

f(x) = x2 + 3x + 2

help me please
1
Expert's answer
2014-06-09T02:02:21-0400
Function could be written as x^2+3x+2=x^2+2x+x+2=x(x+2)+x+2=(x+1)(x+2). From this we conclude that zeros are -1 and -2.

Answer. x=-1 and x=-2.